# Demonstration of sequential processors with quantum advantage and analysis of classical performance limits

### v1

@@ VI Extensions and Outlook
In this paper, we have studied theoretically and experimentally the separation between performance of quantum and classical processors of the same structure tackling the problem of approximation of certain functions of input variables when the variables are available only locally to different modules and the communication between the modules is restricted. In particular, binary functions can be achieved by the quantum sequential processors with qubit or qutrit communication, which can only be approximated by the output of the corresponding classical processors. We showed that existing silicon optics technology allows us to demonstrate this separation.
@@
To prove the separation in the cases tailored to our experiment, we needed to develop a general theoretical framework which allows us to put strict boundaries on the performance of classical processors. We did it two ways. From the properties of specific target binary functions related to the output of our experiment, and from the reduction of the optimization in the set of binary function which some properties to the Ising problem for arbitrary binary targets.
@@
If the target function is an arbitrary binary function, the Ising optimization, or alternatively the optimization of the corresponding correlation function gives the solution to the problem of the best approximation of one arbitrary binary function by another which satisfies some restrictions. In our case, the restriction is to be the output of a sequential processor with communication between the modules restricted to a link with given dimensionality. The restriction has a clear interpretation when we reshape the function in the form of a matrix, the rows of which correspond to the processor response to given outputs of one module. As the number of symbols in the output is restricted, the number of different rows is restricted. The number of different rows in a binary matrix corresponds to the rank. So, the problem of finding the optimal output of a two-module processor with limited communication to an arbitrary binary function directly corresponds to the problem of approximation of an arbitrary binary matrix by a binary matrix of specific rank. The optimality is defined here by the Hamming distance.
@@
Unlike in low rank approximation of real or complex matrices, which can be done by reducing rank in the singular value decomposition, the low rank approximation of binary matrices is known as the hard problem. A side result of our work is the reduction of this problem to the generalized Ising problem, which is also known to be hard in general, but dedicated approximate solvers exist and are still improving.
@@
Another restriction on approximation of the general binary target function of a number of variables can be the output of a multi-module processor. In this case, we approximate a general tensor (indexed by the variables) in terms of a concatenation of tensors of specific ranks. Finding an optimal approximation in terms of the Hamming distance is again a difficult task, which we explicitly reduce to a generalized Ising problem and can attack with existing Ising solvers.
@@
The third related application of the techniques we developed in this paper is the binary matrix completion. We could ask, what is a good fixed rank binary matrix completion of a binary matrix with restricted number of known entries. Again we tacle this problem by creating the binary matrix with highest possible correlation function with the existing part of the reference matrix and with required values of the rank or ranks in different matrix reshapings.
@@
Finally, the direct application of our techniques is the logic and sequential logic design. In fact, given an arbitrary binary target function of a number of variables can be interpreted as the truth table for some logical formulas of the corresponding Boolean variables. The optimal construction of the logical circuits or circuits satisfying specific conditions such as being distributed or sequential is a highly desirable and unsolved problem. Although approximations exist and are widely used, new techniques, such as the one proposed in this paper, can contribute to the design improvement.
@@ Acknowledgements.
We would like to thank Marek Żukowski for valuable discussions. This work was supported by JST Moonshot R&DR\&D, Grant No. JPMJMS226C and Grant No. JPMJMS2061, JST ASPIRE, Grant No. JPMJAP2427, JST COI-NEXT Grant No. JPMJPF2221, JST ERATO Grant No. JPMJER2402, and JSPS KAKENHI Grant Number JP24K00559.

### v2

@@ V Discussion: extension of the results
In this paper, we have studied theoretically and experimentally the separation between performance of quantum and classical processors of the same structure tackling the problem of approximation of certain functions of input variables when the variables are available only locally to different modules and the communication between the modules is restricted. The quantum processor could potentially use intermediate measurements and effectively collect and pass classical communication. Therefore, the quantum processor could simulate a classical one. However, the opposite is not true. In particular, binary functions that can be achieved by the quantum sequential processors with qubit or qutrit communication can only be approximated by the output of the corresponding classical processors. We showed that existing silicon optics technology allows us to demonstrate this separation.
@@
To prove the separation in the cases tailored to our experiment, we needed to develop a general theoretical framework that allows us to put strict boundaries on the performance of classical processors. We did it in two ways. The first one is based on the properties of specific target binary functions related to the output of our experiment. The second way of finding classical bounds does not rely on particular structure of the target function. It comes from finding the minimimum Hamming distance between a given binary function and a set of binary functions that satisfy certain conditions.
@@
In our case, the condition is to be the output of a sequential processor with communication between the modules restricted to a link with given dimensionality. The restriction has a clear interpretation when we reshape the function in the form of a matrix. The rows of the matrix correspond to the processor response to given outputs of one module. As the number of symbols in the output is restricted, the number of different rows is also restricted. The number of different rows in a binary matrix corresponds to its rank. So, the problem of finding the optimal output of a two-module processor with limited communication to an arbitrary binary function directly corresponds to the problem of approximation of an arbitrary binary matrix by a binary matrix of specific rank. The optimality is here defined by the Hamming distance. We reduce this optimization to an Ising problem and solve it in certain cases by existing Ising solvers.
@@
Unlike low-rank approximation of real and complex matrices that can be done by reducing rank in the singular value decomposition, the low-rank approximation of binary matrices is known to be NP hard. A side result of our work is the reduction of this problem to the generalized Ising problem. The general Ising problem is also known to be hard in general, but dedicated approximate solvers exist and are still improving. Notably, quantum computers are candidates for interesting heuristics in the resolution of this problem.
@@
Another restriction on approximation of the general binary target function of a number of variables can be the output of a multimodule processor. In this case, we approximate a general tensor (indexed by variables) in terms of a concatenation of tensors of specific ranks. Finding an optimal approximation, in terms of the Hamming distance, is again a difficult task. We can explicitly reduce it to a generalized Ising problem and approach it with existing Ising solvers.
@@
The third related application of the techniques developed in this paper is the binary matrix completion. We could ask, what is a good fixed rank binary matrix completion of a given binary matrix with restricted number of known entries. Again, we tackle this problem by creating the binary matrix with highest possible correlation function with the existing part of the reference matrix and with constraints on the ranks in different reshapings of the matrix.
@@
Finally, the direct application of our techniques is the logic and sequential logic design. In fact, an arbitrary binary target function of a number of variables can be interpreted as the truth table for some logical formulas of the corresponding Boolean variables. The optimal construction of logical circuits or circuits that meet specific conditions, such as being distributed or sequential, is a highly desirable and unsolved problem. Although approximations exist and are widely used, techniques, such as the one proposed in this paper, can contribute to design improvement.
@@ Acknowledgements.
We thank Marek Żukowski for valuable discussions. This work was supported by JST Moonshot R&DR\&D, Grant No. JPMJMS226C and Grant No. JPMJMS2061, JST ASPIRE, Grant No. JPMJAP2427, JST COI-NEXT Grant No. JPMJPF2221, JST ERATO Grant No. JPMJER2402, and JSPS KAKENHI Grant No. JP24K00559, JST SPRING Grant No. JPMJSP2123.
Appendix A Classical bounds on correlations
In this appendix, we establish bounds on the correlations with the target functions given in Eqs. (8) and (14), which are achievable by outputs of sequential classical processors with one-bit and one-trit communication, respectively, as shown in Figs. 46. The maximum correlation is uniquely defined by the minimum number of errors between the output of the classical processor and the target function. In the following parts, we establish these results as propositions.
A.1 Bound for one-bit processor
Proposition 1.
Proof.
In the proof, we explicitly use the properties of the target function (8). The rows of this matrix correspond to the processor action depending on the first-stage output. As the first module can have only two different symbols, the target reshuffled as a $4\times 16$ matrix needs to be of rank 2. So, two lines need to be identical. Let these lines be first and third (the same can be repeated for all other choices). We add the two rows, obtaining a binary vector of length 16.
We reshape this vector into a matrix $4\times 4$. It needs to be rank 2, as the rows describe the processor’s actions depending on the output from the second stage, which can output at most two symbols. However, we observe that each row is different and that the minimum number of different bits between each pair of rows (minimum Hamming distance) is 2. We need to correct two bits in two rows. In this way, we necessarily introduce four errors.
Finally, we have the remaining lines 2 and 4 of the target reshaped into the matrix $4\times 16$. When we reshape each of these lines into $4\times 4$ matrices, for the same reason as before, we need to correct four additional bits.
Performing simulated annealing combinatorial optimization on Fixstars Amplify we find a strategy for each module of the processor such that the final output approximates the non-zero part of the target with exactly eight errors. The example is shown in Table 1.
This concludes the proof. The minimum eight errors correspond to the maximum value of the correlation between the output and the target equal to 0.25. ∎
Actually, because the size of the problem is relatively small, we could guess the solution with only eight errors. However, we perform the simulation on the annealer to demonstrate that our numerical optimization tool works well.
A.2 Bound for three-module trit processor
Proposition 2.
The number of errors between the output of the three-stage processor with one-trit communication and the nonzero part of the target function (14) (lines 1 to 4) cannot be less than 2.
Proof.
The target function is the first four rows of Eq. (14). As the rows of the matrix correspond to the functions in modules 2 and 3 given one of three symbols output from module 1, the rank of the matrix has to be equal to 3. Because the matrix is binary, the rank restriction requires that two of the rows be identical.
Without loss of generality, assume that the first and third rows are identical (the procedure can be repeated in all other cases). We complete the undefined elements of the first row by the defined elements of the third row. We observe that we have four blocks of size 4. Moreover, all are different, and the number of different bits between each pair of the four blocks is equal to 2. We can reduce the number of these 4-size blocks to three by changing at least two bits. So, the smallest number of changes we would need is 2.
We can give an explicit example of a processor that approximate the target function with only two errors, as shown in Table 2.
A.3 Bound for four-module trit processor
Proposition 3.
Proof.
The proof uses the properties of the target function (14) explicitly.
We reshape the target into the $4\times 64$ matrix. The rows correspond to the processor action depending on the first-stage output. As it can have only three different symbols, the target reshuffled as a $4\times 64$ matrix must be of rank 3. So, two lines need to be identical. Let these lines be the first and second (the same can be repeated for all other choices). We add the two rows, obtaining a binary function of length 64.
We reshape this vector into a matrix $4\times 16$. It needs to be rank 3, as the rows describe the processor’s actions depending on the output from the second stage, which can contain at most three symbols. So, two rows of this matrix have to be the same. However, we observe that the minimum number of different bits between each pair of rows (minimum Hamming distance) is 8. Correcting some entries in one of the rows, we necessarily introduce eight errors in the original target function.
We observe that the remaining lines of the $4\times 16$ matrix contain four different blocks of length 4. To reduce this number to three different blocks of length 4, we need to introduce two additional errors per line. This means that four additional errors are required to correct these parts.
Finally, we have the remaining lines 3 and 4 of the target reshaped into the $4\times 64$ matrix. When we reshape each of these lines into $4\times 16$ matrices, for the same reason as before, we expect that each of them is of rank 3. So, the two lines need to be identical. We observe that when we add two complementary lines we get vectors of length 16 consisting of four blocks of size 4. To reduce this number to 3, we need to introduce two errors per each of the two rows.
Performing simulated annealing combinatorial optimization on Fixstars Amplify, we find an optimal strategy for each module of the processor such that the final output approximates the non-zero part of the target with exactly 16 errors. The example is shown in Table 3.
Appendix B Existence of classical processor with a given output
For simplicity, assume that the number of processing modules is $N=3$. We can solve Task 1. by checking the number of different blocks of size 4, and 16 in the target function. There are 16 blocks of size 4. They are indexed by sequences of bits $(X_{1}X_{2}X_{3}X_{4})$ and defined as
$t^{[4]}(X_{1}\ldots X_{4},X_{5}X_{6})=T(X_{1},\ldots,X_{6}).$
There are four blocks of size 16. They are indexed by sequences of bits $(X_{1}X_{2})$ defined as
$t^{[16]}(X_{1}X_{2},X_{3}\ldots X_{6})=T(X_{1},\ldots,X_{6})$
Theorem 1.
The sequential processor with three stages and one-bit communication can output functions $O(X_{1},...,X_{6})$ with at most two different rows in the matrix $O^{[4]}(X_{1}X_{2},X_{3}...X_{6})$ and at most two different rows in the reshuffled matrix $O^{[16]}(X_{1}...X_{4},X_{5}X_{6})$.
Proof.
There are four possible outputs (one output for each $X_{1}X_{2}$) from the first processor but only two binary symbols 0 or $1$. For each symbol, there is one and only one succession [the row of the matrix $O^{[4]}(X_{1}X_{2},X_{3}...X_{6})$]. This is because the following stage cannot distinguish the same symbols output for different inputs of the first stage. So, there can be only two different rows in matrix $O^{[4]}$. Similarly, there are 16 outputs (one output for each sequence $(X_{1}...X_{4})$) from the second processor, but only two binary symbols. For each symbol, there is a single succession. The row of the matrix $O^{[16]}(X_{1}...X_{4},X_{5}X_{6})$ for a chosen value of $(X_{1}...X_{4})$ is equal to one of these successions. We recognize that the target function consists of only two distinct rows of $O^{[16]}$ and only two distinct rows of $O^{[4]}$. Thus, we can associate symbols 0 or $1$ as the output of the first processor in the positions corresponding to different distinct rows. Similarly, we associate 0 or $1$ as the outputs of the second processor in the positions corresponding to distinct rows of $O^{[16]}$. ∎
The argument can be easily extended to processors with more modules and different limits on communication between them.